Find eigen values and associated eigen vectors of the matrix

A=[ 1 −3 3

−3 −5 3

6 −6 4 ] in the field R. Also find an invertible matrix p such that p−1 Ap is diagonal.


1
Expert's answer
2021-12-24T12:11:40-0500
"A-\lambda I=\begin{bmatrix}
 1-\lambda & -3 & 3 \\
 -3 & -5-\lambda & 3 \\
 6 & -6 & 4-\lambda
\end{bmatrix}"

The characteristic equation

"\det(A-\lambda I)=0"

"(1-\lambda)(-5-\lambda)(4-\lambda)-54+54"

"-18(-5-\lambda)+18(1-\lambda)-9(4-\lambda)=0"

"-\lambda^3+30\lambda+52=0"

"-\lambda^2(\lambda+2)+2\lambda(\lambda+2)+26(\lambda+2)=0"

"\lambda_1=-2"

"\lambda^2-2\lambda-26=0"

"(\lambda-1)^2=27"

"\lambda_2=1-3\sqrt{3}, \lambda_3=1+3\sqrt{3}"

"\lambda_1=-2,\lambda_2=1-3\sqrt{3}, \lambda_3=1+3\sqrt{3}"

These are eigenvalues.


"\lambda=-2"


"\begin{bmatrix}
 1-\lambda & -3 & 3 \\
 -3 & -5-\lambda & 3 \\
 6 & -6 & 4-\lambda
\end{bmatrix}=\begin{bmatrix}
 3 & -3 & 3 \\
 -3 & -3 & 3 \\
 6 & -6 & 6
\end{bmatrix}"

"R_1=R_1\/3"


"\begin{bmatrix}
 1 & -1 & 1 \\
 -3 & -3 & 3 \\
 6 & -6 & 6
\end{bmatrix}"

"R_2=R_2+3R_1"


"\begin{bmatrix}
 1 & -1 & 1 \\
 0 & -6 & 6 \\
 6 & -6 & 6
\end{bmatrix}"

"R_3=R_3-6R_1"


"\begin{bmatrix}
 1 & -1 & 1 \\
 0 & -6 & 6 \\
 0 & 0 & 0
\end{bmatrix}"

"R_2=R_2\/(-6)"


"\begin{bmatrix}
 1 & -1 & 1 \\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{bmatrix}"

"R_1=R_1+R_2"


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{bmatrix}"

If we take "v_3=t," then "v_1=0, v_2=t, v_3=t."

Thus the eigenvector is

"\bold v=\begin{bmatrix}
 0 \\
 1 \\
1
\end{bmatrix}"


"\lambda=1-3\sqrt{3}"

"\begin{bmatrix}
 1-\lambda & -3 & 3 \\
 -3 & -5-\lambda & 3 \\
 6 & -6 & 4-\lambda
\end{bmatrix}"

"=\begin{bmatrix}
 3\sqrt{3} & -3 & 3 \\
 -3 & -6+3\sqrt{3} & 3 \\
 6 & -6 & 3+3\sqrt{3}
\end{bmatrix}"

"R_1=\sqrt{3}R_1/9"

"\begin{bmatrix}
 1 & -\sqrt{3}/3 & \sqrt{3}/3 \\
 -3 & -6+3\sqrt{3} & 3 \\
 6 & -6 & 3+3\sqrt{3}
\end{bmatrix}"

"R_2=R_2+3R_1"

"\begin{bmatrix}
 1 & -\sqrt{3}/3 & \sqrt{3}/3 \\
 0 & -6+2\sqrt{3} & 3+\sqrt{3} \\
 6 & -6 & 3+3\sqrt{3}
\end{bmatrix}"

"R_3=R_3-6R_1"

"\begin{bmatrix}
 1 & -\sqrt{3}/3 & \sqrt{3}/3 \\
 0 & -6+2\sqrt{3} & 3+\sqrt{3} \\
 0 & -6+2\sqrt{3} & 3+\sqrt{3}
\end{bmatrix}"

"R_2=R_2/(-6+2\sqrt{3})"

"\begin{bmatrix}
 1 & -\sqrt{3}/3 & \sqrt{3}/3 \\
 0 & 1 & -1-\sqrt{3}/2 \\
 0 & -6+2\sqrt{3} & 3+\sqrt{3}
\end{bmatrix}"

"R_1=R_1+R_2/\sqrt{3}"

"\begin{bmatrix}
 1 & 0 & -1/2 \\
 0 & 1 & -1-\sqrt{3}/2 \\
 0 & -6+2\sqrt{3} & 3+\sqrt{3}
\end{bmatrix}"

"R_3=R_3+(6-2\sqrt{3})R_2"


"\begin{bmatrix}
 1 & 0 & -1/2 \\
 0 & 1 & -1-\sqrt{3}/2 \\
 0 & 0 & 0
\end{bmatrix}"

If we take "u_3=t," then "u_1=t/2, u_2=(2+\sqrt{3})t/2, u_3=t."

Thus the eigenvector is

"\bold u=\begin{bmatrix}
 1/2 \\
 (2+\sqrt{3})/2 \\
1
\end{bmatrix}"



"\lambda=1+3\sqrt{3}"

"\begin{bmatrix}
 1-\lambda & -3 & 3 \\
 -3 & -5-\lambda & 3 \\
 6 & -6 & 4-\lambda
\end{bmatrix}"

"=\begin{bmatrix}
 -3\sqrt{3} & -3 & 3 \\
 -3 & -6-3\sqrt{3} & 3 \\
 6 & -6 & 3-3\sqrt{3}
\end{bmatrix}"

"R_1=-\sqrt{3}R_1/9"

"\begin{bmatrix}
 1 & \sqrt{3}/3 & -\sqrt{3}/3 \\
 -3 & -6-3\sqrt{3} & 3 \\
 6 & -6 & 3-3\sqrt{3}
\end{bmatrix}"

"R_2=R_2+3R_1"

"\begin{bmatrix}
 1 & \sqrt{3}/3 & -\sqrt{3}/3 \\
 0 & -6-2\sqrt{3} & 3-\sqrt{3} \\
 6 & -6 & 3-3\sqrt{3}
\end{bmatrix}"

"R_3=R_3-6R_1"

"\begin{bmatrix}
 1 & \sqrt{3}/3 & -\sqrt{3}/3 \\
 0 & -6-2\sqrt{3} & 3-\sqrt{3} \\
 0 & -6-2\sqrt{3} & 3-\sqrt{3}
\end{bmatrix}"

"R_2=R_2/(-6-2\sqrt{3})"

"\begin{bmatrix}
 1 & \sqrt{3}/3 & -\sqrt{3}/3 \\
 0 & 1 & -1+\sqrt{3} /2\\
 0 & -6-2\sqrt{3} & 3-\sqrt{3}
\end{bmatrix}"

"R_1=R_1-R_2/\sqrt{3}"

"\begin{bmatrix}
 1 & 0 & -1/2 \\
 0 & 1 & -1+\sqrt{3} /2\\
 0 & -6-2\sqrt{3} & 3-\sqrt{3}
\end{bmatrix}"

"R_3=R_3+(6+2\sqrt{3})R_2"

"\begin{bmatrix}
 1 & 0 & -1/2 \\
 0 & 1 & -1+\sqrt{3} /2\\
 0 & 0 & 0
\end{bmatrix}"


If we take "w_3=t," then "w_1=t/2, w_2=(2-\sqrt{3})t/2, w_3=t."

Thus the eigenvector is

"\bold w=\begin{bmatrix}
 1/2 \\
 (2-\sqrt{3})/2 \\
1
\end{bmatrix}"


Form the matrix "P," whose column "i" is eigenvector no. "i:"


"P=\begin{bmatrix}
 0 & 1/2 & 1/2 \\
 1 & 1+\sqrt{3}/2 & 1-\sqrt{3} /2\\
 1 & 1 & 1
\end{bmatrix}"

Form the matrix "D" whose element at row "i," column "i" is eigenvalue no. "i:"


"D=\begin{bmatrix}
 -2 & 0 & 0 \\
 0 & 1-3\sqrt{3} & 0\\
 0 & 0 & 1+3\sqrt{3}
\end{bmatrix}"

The matrices "P" and "D"  are such that the initial matrix


"A=\begin{bmatrix}
 1 & -3 & 3 \\
 -3 & -5 & 3 \\
 6 & -6 & 4
\end{bmatrix}=PDP^{-1}"